# Conformalized Interactive Imitation Learning: Handling Expert Shift & Intermittent Feedback

### v1

@@ 1 Introduction
Figure 1: Conformalized Interactive Imitation Learning. The robot learns an initial policy via imitation learning to wipe a line drawing by following a straight path. When the initial policy is deployed, ConformalDAgger calibrates the robot’s uncertainty (represented by red prediction interval boxes) based on feedback received from the human during the interactive IL loop. (left) The robot refrains from asking questions when its uncertainty is low because the expert policy at deployment time is aligned to the demonstrations on which the learner was trained. (right) When the human shifts in their task strategy from the training data distribution, the robot uncertainty increases and enables it to actively query for more labels.
@@ 2 Related Work
Interactive IL is a branch of imitation learning wherein a robot learner can query a (human) expert to receive additional labels either during or after task execution (Celemin et al., 2022). A foundational approach for interactive IL is DAgger (Dataset Aggregation) (Ross et al., 2011), which iteratively augments the training dataset by aggregating data from the expert and learner policies and assuming the expert is stationary. In the online case, the robot learner can cede control to the expert at any time to get additional state-action data (also known as robot-gated feedback) or the expert can actively intervene at any time (also known as human-gated feedback) (Kelly et al., 2019). From the learner’s perspective, a key question is when to request feedback from the expert so that it minimizes expert effort but also minimizes negative events caused by an erroneous policy (e.g., running into a wall). On one hand, prior works focus on minimizing human effort by constraining robot requests via a limited human attention model (Hoque et al., 2023) or a budget of human interventions (Hoque et al., 2021a; b). On the other hand, prior works prioritize deployment-time safety by classifying safe versus unsafe states (SafeDAgger (Zhang & Cho, 2017), Replay Estimation (Swamy et al., 2022)), estimating uncertainty via ensemble disagreement (EnsembleDAgger (Menda et al., 2019)), or Monte Carlo dropout (Cui et al., 2019). We present ConformalDAgger, which uses our novel uncertainty quantification method grounded in conformal prediction to adaptively increase the learner requests for help under uncertainty and decrease the number of requests when confident.
Interacting with Non-stationary Experts. A core kind of distribution shift we study in this work is human expert distribution shift. This can occur for a variety of reasons, from human teleoperators having biases when teaching robots (Thomaz et al., 2006; Thomaz & Breazeal, 2008), to generating different data distributions because of suboptimality (Grüne-Yanoff, 2015; Thompson, 1999) or varied risk tolerance (Kwon et al., 2020). Moreover, simply observing robot behavior can influence how humans react or teach robots over time (Hong et al., 2024; Sagheb et al., 2023; Xie et al., 2021).
Despite its prevalence in the real world, the majority of interactive IL works assume a stationary expert policy (Likmeta et al., 2021; Shin et al., 2023; Zheng et al., 2022). Our conformalized imitation learning approach takes a key step towards closing this gap by developing an online conformal prediction algorithm that can account for expert distribution shift.
@@
Online Conformal Prediction. Conformal prediction is a distribution-free uncertainty quantification method for constructing prediction intervals for both classification and regression problems (Angelopoulos & Bates, 2023; Romano et al., 2019; 2020; Zaffran et al., 2022), as well as for offline or online data. We focus on the online setting (e.g., timeseries) where uncertainty quantification is performed on streaming pairs of input-label data that are not necessarily i.i.d. (Gibbs & Candes, 2021). Broadly speaking, there are two predominant algorithms in this setting: adaptive conformal inference (ACI) (Gibbs & Candes, 2021; Gibbs & Candès, 2024; Bhatnagar et al., 2023; Zaffran et al., 2022) and quantile tracking (QT) (Angelopoulos et al., 2024a). Both are online gradient descent-based methods which guarantee asymptotic coverage in the online setting. Unlike ACI which is prone to infinitely sized intervals after a series of miscoverage events, QT directly estimates the value of the empirical quantile itself, ensuring coverage with finite intervals. In this work, we relax the assumption that labels must be observed at each time point in the streaming data and extend the online conformal paradigm to ensure coverage in the intermittent label regime.
@@ 3 Online Conformal Prediction with Intermittent Labels
From the uncertainty quantification side, our core technical contribution is extending online conformal prediction to settings where ground truth labels are intermittently observed. We present our algorithm in the context of quantile tracking for relevance to our interactive IL experiments. However, we also derive an extension of ACI (Gibbs & Candes, 2021) to intermittent labels in the Appendix Section B.
@@
Background: Quantile Tracking. The quantile tracking (QT) algorithm from Angelopoulos et al. (2024b) implicitly seeks to track the value of the $1-\alpha$ quantile via online gradient descent on the quantile loss (Koenker & Bassett Jr, 1978).
As in Angelopoulos et al. (2024b), we leverage a bounded nonconformity score function: $s:\mathcal{X}\times\mathcal{Y}\rightarrow[0,B]$ where $0<B<\infty$, to quantify the error made by the initial prediction of the base model. We assume the nonconformity score function $s(x_{t},y_{t})$ is negatively oriented (lower values indicate less nonconformity or greater prediction accuracy). Let $q_{t}$ represents the estimated $1-\alpha$ quantile of the score sequence $s_{t},t\in\mathbb{N}$. Prediction intervals are constructed using the nonconformity score function:
@@
Intuitively, this update increases the quantile threshold when the model continuously miscovers and decreases the quantile when the model coverage is performant. For arbitrary step size $\gamma_{t}$ with no assumptions on the sequence of data points $(x_{1},y_{1}),(x_{2},y_{2}),...$ and an initial quantile threshold $q_{1}\in[0,B]$, the quantile tracking update satisfies Equation 3 (Theorem 2 of Angelopoulos et al. (2024b)):
@@
Refer to full proof in Appendix Section A. Note that when feedback is constant (i.e. $p_{t}=1$), $\text{obs}_{t}=1$ for all timesteps, and IQT reduces to online quantile tracking with arbitrary step sizes.
Practical Note: Choosing Gamma. In practice, an important decision is the choice of $\gamma_{t}$. Prior works in quantile tracking with constant labels (Angelopoulos et al., 2024a) choose $\gamma_{t}=\text{lr}\hat{B}_{t}$, where $\hat{B}_{t}:=\max_{t-k\leq r\leq t-1}s(x_{r},y_{r})$, $k$ represents a look-back window across the previous timesteps with observed ground truth labels, and lr is a small constant. On our domains, we find that choosing $\gamma_{t}$ such that it contains $\hat{B}_{t}$ improves empirical coverage over the choice of a constant $\gamma_{t}=\gamma$.
@@ 3.1 Experiments: Intermittent Quantile Tracking on Standard Timeseries Data
Figure 3: IQT on Amazon Stocks with AR base model: Coverage & Interval Visualization. We set $p_{t}=0.1,\forall t$ to simulate seeing the true price only 10% of the time and show the prediction interval for 1 seed and coverage averaged over 5 seeds (shaded error on coverage plots show standard deviation). IQT-pd is prone to large intervals when lr is large, and boosts coverage over IQT-pi when lr is small.
@@
Setup. We test on three benchmark datasets from Angelopoulos et al. (2024a): (1) Amazon stock prices, (2) Google stock prices (Nguyen, 2018), and the (3) Elec2 dataset (Harries, 1999). We test four base prediction models, $\hat{f}$, all trained via darts (Herzen et al., 2022) to see if IQT consistently maintains coverage close to the desired level. We present the Autoregressive (AR) model with 3 lags for brevity in the main text and defer the other model results to the Appendix. Our nonconformity score is the asymmetric (signed) residual score. We measure (1) marginal coverage over the time series, (2) longest miscoverage error sequence, and (3) mean prediction interval size.
Observation Models. We test three different observation frequencies: (1) infrequent $p_{t}=0.1,\forall t$, (2) partial $p_{t}=0.5,\forall t$, and (3) frequent $p_{t}=0.9,\forall t$. We focus in this section on the results for infrequent observations, and discuss further results in the Appendix Section C.
@@
Results. We center our discussion on the Amazon stock price results for IQT-pd and IQT-pi with AR base model under ground truth labels observed with $p_{t}=0.1$ frequency, as the intermittent setting is our focus. See Appendix Section C for further results on other base models and datasets. Inversely scaling the quantile tracking step size by $p_{t}$ causes IQT-pd to construct larger intervals than IQT-pi. Figure 3 shows the prediction coverage as a moving average (window=50) for 5 random seeds and interval sizes for one seed. When lr is high (1.0), IQT-pi and IQT-pd achieve comparable coverage, but IQT-pd is prone to constructing much larger prediction intervals than IQT-pi. Smaller and midsized learning rates ($\text{lr}=0.01,0.1$) regulate the size of IQT-pd intervals, leading to tight prediction intervals which maintain desired coverage levels.
@@ 4 ConformalDAgger: A Calibrated Approach to Asking for Expert Feedback
Observation Model.
A key component of IQT is the observation likelihood model, $p_{t}=\mathbb{P}(\text{obs}_{t}=1\mid x_{t})$. A nice byproduct of this model is that we can naturally derive a feedback model that is simultaneously human- and robot-gated by decomposing the observation model into the combination of both gating functions.
Intuitively, the likelihood of observing the expert feedback at $t$ is given by the probability that the human chooses to give feedback or the robot asks the expert for feedback. Let $\text{obs}^{h}_{t}\in\{0,1\}$ be a random variable representing observing a human-gated feedback (if $\text{obs}^{h}_{t}=1$) where the expert initiates providing an action label $a^{h}_{t}=\pi^{h}(x_{t})$ for $x_{t}$. Let $\text{obs}^{r}_{t}\in\{0,1\}$ be a random variable representing robot-gated feedback, where the robot asks the expert for a label if $\text{obs}^{r}_{t}=1$. We assume that if the robot elects to ask a question, the human will respond with probability 1. Our observation model takes the form:
$p_{t}$ $:=\mathbb{P}(\text{obs}_{t}=1\mid x_{t})=\mathbb{P}(\text{obs}^{h}_{t}=1\lor\text{obs}^{r}_{t}=1\mid x_{t})$ (7)
$=\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})+\mathbb{P}(\text{obs}^{r}_{t}=1\mid x_{t})-\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})\mathbb{P}(\text{obs}^{r}_{t}=1\mid x_{t})$ (8)
While our approach is not prescriptive about these models, in our experiments we model $\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})=c$ as a small constant to represent infrequent human interventions. We describe the robot-gated likelihood model $\mathbb{P}(\text{obs}^{r}_{t}=1\mid x_{t})$ below, informed by our uncertainty estimates.
@@
7:         if robot queries (sample $\mathbb{P}(\text{obs}^{r}_{t}\mid o_{t})$) or human intervenes (sample $\mathbb{P}(\text{obs}^{h}_{t}\mid o_{t})$then
8:              IQT update: $q^{lo}_{t+1}\leftarrow q^{lo}_{t}+\frac{\gamma_{t}}{p_{t}}(\text{err}_{t}-\alpha)\text{obs}_{t}$  with expert action $a^{h}_{t}\leftarrow\pi^{h}_{i}(x_{t})$.
9:              IQT update: $q^{hi}_{t+1}\leftarrow q^{hi}_{t}+\frac{\gamma_{t}}{p_{t}}(\text{err}_{t}-\alpha)\text{obs}_{t}$  with expert action $a^{h}_{t}\leftarrow\pi^{h}_{i}(x_{t})$.
@@
11:              IQT update: $q^{lo}_{t+1}\leftarrow q^{lo}_{t}$.
12:              IQT update: $q^{hi}_{t+1}\leftarrow q^{hi}_{t}$.
13:     Aggregated dataset with observed state and expert action pairs: $\mathcal{D}_{i+1}\leftarrow\mathcal{D}_{i}\cup\{(x,\pi^{h}_{i}(x))\}$
14:     Retrain learner: $\pi^{r}_{i+1}\leftarrow\operatorname*{arg\,min}_{\pi}\mathcal{L}(\mathcal{D}_{i+1})$
Quantifying Uncertainty: IQT.
In the interactive IL setting, IQT begins with initial upper and lower quantiles in the action (i.e. “label”) space, $q^{lo}_{0},q^{hi}_{0}\in\mathcal{A}$. The nonconformity score is a residual on the predicted ($a^{r}$) versus expert ($a^{h}$) action. Let $s^{lo}_{t}(a^{r}_{t},a^{h}_{t})=a^{r}_{t}-a^{h}_{t}$ be the lower residual (referred to as $s^{lo}_{t}$ for brevity), and $s^{hi}_{t}(a^{r}_{t},a^{h}_{t})=a^{h}_{t}-a^{r}_{t}$ be the upper residual (referred to as $s^{hi}_{t}$). If the expert action $a^{h}_{t}$ is observed, we compute the upper and lower miscoverage of $q^{lo}_{t},q^{hi}_{t}$ as two indicator vectors: $\text{err}_{t}^{lo}=(s^{lo}_{t}<q^{lo}_{t})$ and $\text{err}_{t}^{hi}=(s^{hi}_{t}<q^{hi}_{t})$. IQT then updates the quantile estimates online to obtain $q^{lo}_{t+1},q^{hi}_{t+1}$ via the update rule from  equation 5. At the next timestep, the adjusted prediction interval is constructed with $C_{t+1}(x_{t+1})=[a^{r}_{t+1}-q^{lo}_{t+1},a^{r}_{t+1}+q^{hi}_{t+1}]$. If the expert action is not observed, $\text{obs}_{t}=0$, then $q^{lo}_{t+1}=q^{lo}_{t}$ and $q^{hi}_{t+1}=q^{hi}_{t}$ and the prediction interval size remains the same. Note that although $\text{err}_{t}$ is not known in the case where the expert does not provide an action label, IQT does not require it; IQT simply makes no change to the quantile estimate. In our simulated and hardware experiments, our action space, $q^{lo}_{0}$, and $q^{hi}_{0}$, are vector-valued. Our experiments instantiate IQT for continuous vectors, but the approach extends to discrete-valued action spaces.
Leveraging Uncertainty: Asking for Help. Finally, the calibrated intervals $C_{t+1}(x_{t+1})$ constructed by IQT enable us to design a new robot-gated feedback mechanism. Specifically, the model $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})$ is informed by the calibrated interval size, $u(x_{t+1}):=\mid\mid C_{t+1}(x_{t+1})\mid\mid_{2}$. In our simulated experiments, we use $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})=\sigma(\beta[u(x_{t};\pi^{r})-\tau])$ where $\tau$ acts as some uncertainty threshold on $u(x_{t};\pi^{r})$, $\beta$ is a temperature hyperparameter, and $\sigma$ is the sigmoid function. In our hardware experiments, we implement $\tau$ to as a hard threshold on on $u(x_{t};\pi^{r})$ above which $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})=1$, and below which $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})=0$.
@@
We note that two features of our setting make it difficult to prove we maintain DAgger performance guarantees: intermittent expert feedback even with a stationary expert, and the possibility of a time-varying expert even with feedback at every state along learner rollouts. Nevertheless, similar to Kelly et al. (2019) and Menda et al. (2019), our approach offers a practical approach to contend with these real-world challenges.
5 Simulated Interative Imitation Learning Experiments

### v2

@@ 1 Introduction
Figure 1: Conformalized Interactive Imitation Learning. The robot learns an initial policy via imitation learning to wipe a line drawing by following a straight path. When the initial policy is deployed, ConformalDAgger calibrates the robot’s uncertainty (represented by red prediction interval boxes) based on feedback received from the human during the interactive IL loop. (left) The robot refrains from asking questions when its uncertainty is low because the expert policy at deployment time is aligned to the demonstrations on which the learner was trained. (right) When the human shifts in their task strategy, uncertainty increases and the robot starts to query for more expert labels.
@@ 2 Related Work
Interactive IL is a branch of imitation learning wherein a robot learner can query a (human) expert to receive additional labels either during or after task execution (Celemin et al., 2022). The utility of in-the-loop expert interventions is mitigating distribution shift between training and deployment of the policy. Distribution shift can stem from a multitude of reasons, including covariate shift (Spencer et al., 2021) (e.g., encountering new observations of the environment during policy execution), or expert shift: wherein the expert’s latent strategy changes over time (Hong et al., 2024; Sagheb et al., 2023; Xie et al., 2021) A foundational approach for interactive IL is DAgger (Dataset Aggregation) (Ross et al., 2011), which iteratively augments the training dataset by aggregating data from the expert and learner policies and assuming the expert is stationary. In the online case, the robot learner can cede control to the expert at any time to get additional state-action data (also known as robot-gated feedback) or the expert can actively intervene at any time (also known as human-gated feedback) (Kelly et al., 2019). From the learner’s perspective, a key question is when to request feedback from the expert so that it minimizes expert effort but also minimizes negative events caused by an erroneous policy (e.g., running into a wall). On one hand, prior works focus on minimizing human effort by constraining robot requests via a limited human attention model (Hoque et al., 2023) or a budget of human interventions (Hoque et al., 2021a; b). On the other hand, prior works prioritize deployment-time safety by classifying safe versus unsafe states (SafeDAgger (Zhang & Cho, 2017), Replay Estimation (Swamy et al., 2022)), estimating uncertainty via ensemble disagreement (EnsembleDAgger (Menda et al., 2019)), or Monte Carlo dropout (Cui et al., 2019). We present ConformalDAgger, which uses our novel uncertainty quantification method grounded in conformal prediction to adaptively increase the learner requests for help under uncertainty and decrease the number of requests when confident.
@@
Online Conformal Prediction. Conformal prediction is a distribution-free uncertainty quantification method for constructing prediction intervals for both classification and regression problems (Angelopoulos & Bates, 2023; Romano et al., 2019; 2020; Zaffran et al., 2022), as well as for offline or online data. We focus on the online setting (e.g., timeseries) where uncertainty quantification is performed on streaming pairs of input-label data that are not necessarily i.i.d. (Gibbs & Candes, 2021). Broadly speaking, there are two predominant algorithms in this setting: adaptive conformal inference (ACI) (Gibbs & Candes, 2021; Gibbs & Candès, 2024; Bhatnagar et al., 2023; Zaffran et al., 2022) and quantile tracking (QT) (Angelopoulos et al., 2023). Both are online gradient descent-based methods which guarantee asymptotic coverage in the online setting. Unlike ACI which is prone to infinitely sized intervals after a series of miscoverage events, QT directly estimates the value of the empirical quantile itself, ensuring coverage with finite intervals. In this work, we relax the assumption that labels must be observed at each time point in the streaming data and extend the online conformal paradigm to ensure coverage in the intermittent label regime.
@@ 3 Online Conformal Prediction with Intermittent Labels
From the uncertainty quantification side, our core technical contribution is extending online conformal prediction to settings where ground truth labels are intermittently observed. We present our algorithm in the context of quantile tracking for relevance to our interactive IL experiments. However, we also derive an extension of ACI (Gibbs & Candes, 2021) to intermittent labels in the Appendix Section D.
@@
Background: Quantile Tracking. The quantile tracking (QT) algorithm from Angelopoulos et al. (2024) implicitly seeks to track the value of the $1-\alpha$ quantile via online gradient descent on the quantile loss (Koenker & Bassett Jr, 1978). As in Angelopoulos et al. (2024), we leverage a bounded nonconformity score function: $s:\mathcal{X}\times\mathcal{Y}\rightarrow[0,B]$ where $0<B<\infty$, to quantify the error made by the initial prediction of the base model. We assume the nonconformity score function $s(x_{t},y_{t})$ is negatively oriented (lower values indicate less nonconformity or greater prediction accuracy). Let $q_{t}$ represents the estimated $1-\alpha$ quantile of the score sequence $s_{t},t\in\mathbb{N}$. Prediction intervals are constructed using the nonconformity score function:
@@
Intuitively, this update increases the quantile threshold when the model continuously miscovers and decreases it when the model coverage is performant. For arbitrary step size $\gamma_{t}$ with no assumptions on the sequence of data points $(x_{1},y_{1}),(x_{2},y_{2}),...$ and an initial quantile threshold $q_{1}\in[0,B]$, the quantile tracking update satisfies Equation 3 (Theorem 2 of Angelopoulos et al. (2024)):
@@
This guarantees IQT gives the $1-\alpha$ long-term empirical coverage frequency, even with intermittent ground truth access, regardless of the underlying data generation process. While these asymptotic coverage guarantees are sound in theory, we believe it is important to acknowledge that typical robot deployment conditions are finite-horizon. Nevertheless our empirical findings in Section 5 indicate the practical utility of IQT’s uncertainty estimates to obtain extra human feedback during deployment. Refer to full proof in Appendix Section C. Note that when feedback is constant (i.e. $p_{t}=1$), $\text{obs}_{t}=1$ for all timesteps, and IQT reduces to online quantile tracking with arbitrary step sizes.
Practical Note: Choosing Gamma. In practice, an important decision is the choice of $\gamma_{t}$. Prior works in quantile tracking with constant labels (Angelopoulos et al., 2023) choose $\gamma_{t}=\text{lr}\hat{B}_{t}$, where $\hat{B}_{t}:=\max_{t-k\leq r\leq t-1}s(x_{r},y_{r})$, $k$ represents a look-back window across the previous timesteps with observed ground truth labels, and lr is a small constant. On our domains, we find that choosing $\gamma_{t}$ such that it contains $\hat{B}_{t}$ improves empirical coverage over the choice of a constant $\gamma_{t}=\gamma$.
@@ 3.1 Experiments: Intermittent Quantile Tracking on Standard Timeseries Data
Figure 3: IQT on Amazon Stocks with AR base model: Coverage & Interval Visualization. We set $p_{t}=0.1,\forall t$ to see true price labels only 10% of the time and show the prediction interval for 1 seed and coverage averaged over 5 seeds (shaded error on coverage plots show std deviation). IQT-pd is prone to large intervals when lr is large, and boosts coverage over IQT-pi when lr is small.
@@
Setup. We test on three benchmark datasets from Angelopoulos et al. (2023): (1) Amazon stock prices, (2) Google stock prices (Nguyen, 2018), and the (3) Elec2 dataset (Harries, 1999). We test four base prediction models, $\hat{f}$, all trained via darts (Herzen et al., 2022) to see if IQT consistently maintains coverage close to the desired level. We present the Autoregressive (AR) model with 3 lags for brevity in the main text and defer the other model results to the Appendix. Our nonconformity score is the asymmetric (signed) residual score. We measure (1) marginal coverage over the time series, (2) longest miscoverage error sequence, and (3) mean prediction interval size.
Observation Models. We test three different observation frequencies: (1) infrequent $p_{t}=0.1,\forall t$, (2) partial $p_{t}=0.5,\forall t$, and (3) frequent $p_{t}=0.9,\forall t$. We focus in this section on the results for infrequent observations, and discuss further results in the Appendix Section E.
@@
Results. We center our discussion on the Amazon stock price results for IQT-pd and IQT-pi with AR base model under ground truth labels observed with $p_{t}=0.1$ frequency, as the intermittent setting is our focus. See Appendix Section E for further results on other base models and datasets. Inversely scaling the quantile tracking step size by $p_{t}$ causes IQT-pd to construct larger intervals than IQT-pi. Figure 3 shows the prediction coverage as a moving average (window=50) for 5 random seeds and interval sizes for one seed. When lr is high (1.0), IQT-pi and IQT-pd achieve comparable coverage, but IQT-pd is prone to constructing much larger prediction intervals than IQT-pi. Smaller and midsized learning rates ($\text{lr}=0.01,0.1$) regulate the size of IQT-pd intervals, leading to tight prediction intervals which maintain desired coverage levels.
@@ 4 ConformalDAgger: A Calibrated Approach to Asking for Expert Feedback
Observation Model. A key component of IQT is the observation likelihood model, $p_{t}=\mathbb{P}(\text{obs}_{t}=1\mid x_{t})$. A nice byproduct of this model is that we can naturally derive a feedback model that is simultaneously human- and robot-gated by decomposing the observation model into the combination of both gating functions. Intuitively, the likelihood of observing the expert feedback at $t$ is given by the probability that the human chooses to give feedback or the robot asks the expert for feedback. Let $\text{obs}^{h}_{t}\in\{0,1\}$ be a random variable representing observing a human-gated feedback (if $\text{obs}^{h}_{t}=1$) where the expert initiates providing an action label $a^{h}_{t}=\pi^{h}(x_{t})$ for $x_{t}$. Let $\text{obs}^{r}_{t}\in\{0,1\}$ be a random variable representing robot-gated feedback, where the robot asks the expert for a label if $\text{obs}^{r}_{t}=1$. We assume that if the robot elects to ask a question, the human will respond with probability 1. Our observation model takes the form: $p_{t}:=\mathbb{P}(\text{obs}_{t}=1\mid x_{t})=\mathbb{P}(\text{obs}^{h}_{t}=1\lor\text{obs}^{r}_{t}=1\mid x_{t})=\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})+\mathbb{P}(\text{obs}^{r}_{t}=1\mid x_{t})-\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})\mathbb{P}(\text{obs}^{r}_{t}=1\mid x_{t}).$ While our approach is not prescriptive about these models, in our experiments we model $\mathbb{P}(\text{obs}^{h}_{t}=1\mid x_{t})=c$ as a small constant to represent infrequent human interventions. We describe the robot-gated likelihood model $\mathbb{P}(\text{obs}^{r}_{t}=1\mid x_{t})$ below, informed by our uncertainty estimates.
@@
7:         Compute robot query likelihood (e.g. $\mathbb{P}(\text{obs}^{r}_{t}\mid o_{t};\tau)=1$ if $u(x_{t};\pi^{r})>\tau$, else 0.)
8:         if robot queries (w.p. $\mathbb{P}(\text{obs}^{r}_{t}\mid o_{t};\tau)$) or human intervenes (w.p. $\mathbb{P}(\text{obs}^{h}_{t}\mid o_{t})$then
9:              IQT update: $q^{lo,hi}_{t+1}\leftarrow q^{lo,hi}_{t}+\frac{\gamma_{t}}{p_{t}}(\text{err}_{t}-\alpha)\text{obs}_{t}$  w/ expert action $a^{h}_{t}\leftarrow\pi^{h}_{i}(x_{t})$.
@@
11:              IQT update: $q^{lo,hi}_{t+1}\leftarrow q^{lo,hi}_{t}$
12:     Aggregated dataset with observed state and expert action pairs: $\mathcal{D}_{i+1}\leftarrow\mathcal{D}_{i}\cup\{(x,\pi^{h}_{i}(x))\}$
13:     Retrain learner: $\pi^{r}_{i+1}\leftarrow\operatorname*{arg\,min}_{\pi}\mathcal{L}(\mathcal{D}_{i+1})$
Quantifying Uncertainty: IQT. In the interactive IL setting, IQT begins with initial upper and lower quantiles in the action (i.e. “label”) space, $q^{lo}_{0},q^{hi}_{0}\in\mathcal{A}$. The nonconformity score is a residual on the predicted ($a^{r}$) versus expert ($a^{h}$) action. Let $s^{lo}_{t}(a^{r}_{t},a^{h}_{t})=a^{r}_{t}-a^{h}_{t}$ be the lower residual (referred to as $s^{lo}_{t}$ for brevity), and $s^{hi}_{t}(a^{r}_{t},a^{h}_{t})=a^{h}_{t}-a^{r}_{t}$ be the upper residual (referred to as $s^{hi}_{t}$). If the expert action $a^{h}_{t}$ is observed, we compute the upper and lower miscoverage of $q^{lo}_{t},q^{hi}_{t}$ as two indicator vectors: $\text{err}_{t}^{lo}=(s^{lo}_{t}<q^{lo}_{t})$ and $\text{err}_{t}^{hi}=(s^{hi}_{t}<q^{hi}_{t})$. IQT then updates the quantile estimates online to obtain $q^{lo}_{t+1},q^{hi}_{t+1}$ via the update rule from  equation 5. At the next timestep, the adjusted prediction interval is constructed with $C_{t+1}(x_{t+1})=[a^{r}_{t+1}-q^{lo}_{t+1},a^{r}_{t+1}+q^{hi}_{t+1}]$. If the expert action is not observed, $\text{obs}_{t}=0$, then $q^{lo}_{t+1}=q^{lo}_{t}$ and $q^{hi}_{t+1}=q^{hi}_{t}$ and the prediction interval size remains the same. Note that although $\text{err}_{t}$ is not known in the case where the expert does not provide an action label, IQT does not require it; IQT simply makes no change to the quantile estimate. In our simulated and hardware experiments, our action space, $q^{lo}_{0}$, and $q^{hi}_{0}$, are vector-valued. Our experiments instantiate IQT for continuous vectors, but the approach extends to discrete-valued action spaces.
Leveraging Uncertainty: Asking for Help. Finally, the calibrated intervals $C_{t+1}(x_{t+1})$ constructed by IQT enable us to design a new robot-gated feedback mechanism. Specifically, the model $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})$ is informed by the calibrated interval size, $u(x_{t+1}):=||C_{t+1}(x_{t+1})||_{2}$. In simulation, we use $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})=\sigma(\beta[u(x_{t};\pi^{r})-\tau])$ where $\tau$ acts as an uncertainty threshold on $u(x_{t};\pi^{r})$, $\beta$ is a temperature hyperparameter, and $\sigma$ is the sigmoid function. In hardware experiments, we set $\tau$ as a hard threshold on on $u(x_{t};\pi^{r})$ above which $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})=1$, and below which $\mathbb{P}(\text{obs}^{r}_{t}\mid x_{t})=0$.
@@
5 Simulated Interactive Imitation Learning Experiments